OpenStudy (anonymous):

2. Twenty people entered the 5000 meter swim at a local pool. a. How many ways can places 1st – 6th be awarded? b. How many ways can the top 8 be determined

OpenStudy (mathstudent55):

This is a question in which the order matters. Does that make it a combination or permutation problem?

OpenStudy (anonymous):

permutation

OpenStudy (mathstudent55):

Correct. You need the formula for choosing r items out of n items. \(_nP_r = \dfrac{n!}{(n - r)!}\)

OpenStudy (mathstudent55):

In your case, for part a, n = 20, and r = 6. For part b, n = 20, and r = 8.

OpenStudy (anonymous):

okay for part A: 20P6 = 20/(20-6) = 20/14 = 20P8 = 20/(20-8) = 20/12 = do you think you could help me understand what to do from here? thank you so much for your help by the way

OpenStudy (anonymous):

the second equation is for part b sorry

OpenStudy (mathstudent55):

Don't forget the factorial signs.

OpenStudy (mathstudent55):